Pam is playing with red and black marbles. The number of red marbles she has is three more than twice the number of black marbles she has. She has 42 in all. How many red marbles does Pam have?
Answer by checkley77(12569)

R=2B+3
R+B=42
(2B+3)+B=42
3B=42-3
3B=39
B=39/3
B=13 BLACK MARBLES.
R=2*13+3
R=26+3
R=29 RED MARBLES.
PROOF:
13+29=42
42=42